OpenStudy (anonymous):

Describe the transformations on the sine curve to get the graph of y = 3sin(4x + π ).

OpenStudy (perl):

here you can reverse engineer it. we know the form Y = A sin (2pi / per. ( x - p.s.)) + v.s.

OpenStudy (perl):

so one way to do this is to rewrite the given sine function and turn it into the form we know about, using algebra y = 3sin(4x + pi) y = 3sin(4x + pi ) + 0 y = 3 sin ( 4( x + pi/4) ) + 0 y = 3 sin ( 4 ( x - (-pi/4 ) ) + 0 y = 3 sin ( 2pi / (pi/2) ( x - (-pi/4) ) ) + 0

OpenStudy (perl):

now it is in the form Y = A sin (2pi / per. ( x - p.s.)) + v.s. also note that A could be negative, in which case |A| is amplitude